# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ 0.B.3 Undecidability of the Inclusion Problem for Several Basic Relations
with $(x_{1},x_{1}^{\prime}),(x_{2_{1}},x_{2_{2}}),(x_{2_{1}},x_{2_{3}}),(x_{2_{1}},x_{2_{4}}),(x_{2_{1}},x_{2_{5}}),(x_{i_{1}},x_{i_{2}})\in R$ for all $i\in\{3,\ldots,n\}$ and there is a morphism $\sigma:X^{*}\rightarrow\Sigma^{*}$ with
$\sigma(\alpha)=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$, then $\sigma(x_{1})=\sigma(x_{1}^{\prime})=a_{1}$ and $\sigma(x_{i_{j}})=a_{i}$ for each $i\in\{1,\ldots,n\}$.
Proof
@@
must hold. Now, $\sigma(x_{2_{j}})=a_{2}$ must hold for every $j\in\{1,2,3,4,5\}$ since
$(x_{2_{1}},x_{2_{2}}),(x_{2_{1}},x_{2_{3}}),(x_{2_{1}},x_{2_{4}}),(x_{2_{1}},x_{2_{5}}),(x_{i_{1}},x_{i_{2}})\in R$ for all $i\in\{3,\ldots,n\}$. Thus, we also get that $\sigma(x_{i_{1}})=\sigma(x_{i_{2}})=a_{i}$ for every $i\in\{3,\ldots,n\}$.
@@
for each $i\in[n]$. Again, all $y_{j,k},\hat{y}_{j,k},\ddot{z}_{j,k}^{\prime}$ are pairwise different variables with $(\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i-1,i}^{\prime}),(\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i,i}^{\prime})\in R$. Now Lemma˜10 applies (mutatis mutandis) as for binary alphabets, and since all substitutions of good form behave for $\Sigma$ as for the binary alphabet, we can use the very same predicates and the same reasoning as before to prove undecidability of the inclusion problem of Theorem˜4.1.
This concludes the proof of Theorem˜4.1.
Appendix 0.C Additional Content for Equivalence Problem
Proof of Lemma˜4:
Proof
Assume $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$. First of all, we know $|\alpha|=|\beta|$ as otherwise, if w.l.o.g. $|\alpha|>|\beta|$, then no shortest word $w\in L_{NE}(\beta,[r_{\beta}])$ obtained by substitutions $h$ where variables are replaced with letters can be obtained from $(\alpha,r_{\alpha})$. So, from now on, assume $|\alpha|=|\beta|$. Suppose $\alpha\neq\beta$. Then there exists a first position $i\in[|\alpha|]$ such that $\alpha[i]\neq\beta[i]$. First, if $\alpha[i],\beta[i]\in\Sigma$ are both terminal letters, then using a substitution $h$ that replaces all variables with letters in $\alpha$ and $\beta$ always results in words $h(\alpha)$ and $h(\beta)$ for which $h(\alpha)[i]\neq h(\beta)[i]$, a contradiction. So, assume w.l.o.g. that $\alpha[i]\in\Sigma$ is a terminal letter and $\beta[i]\in X$ is a variable. Then, again, any substitution $h\in H_{r_{\beta}}$ that replaces all variables in $\beta$ with letters and sets $h(\beta[i])$ to a letter which is not $\alpha[i]$ produces a word that cannot be in $L_{NE}(\alpha,[r_{\alpha}])$, another contradiction to $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$. As we assume all variables to be introduced in an increasing manner, $\alpha=\beta$. ∎
Proof of Lemma˜5:
Proof
($\Leftarrow$): Assume $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$. Then $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$. So, we trivially get $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$.
($\Rightarrow$) Assume $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$. By Lemma 4 we have $\alpha=\beta$. The relational constraints $r_{\alpha}]$ and $r_{\beta}]$ imply blocks of equivalently substituted variables under $\tilde{r}$. Suppose $r_{\alpha}]\neq[r_{\beta}]$. Let $\alpha=\alpha_{p}x_{i}\alpha_{s}=\beta$ for some $\alpha_{p},\alpha_{s}\in\mathtt{RelPat}_{\Sigma,R}$ such that $x_{i}$ is the first variable for which there is a difference in the relations, i.e., we get that w.l.o.g. $x_{i}\in\operatorname{var}(\alpha)$ is related to some other variable $x_{j}\in\operatorname{var}(\alpha)$ in $r_{\alpha}]$ but not in $r_{\beta}]$ (note that $\operatorname{var}(\alpha)=\operatorname{var}(\beta)$). So, there exists $(\tilde{r},x_{i},x_{j})\in[r_{\alpha}]$ such that $(\tilde{r},x_{i},x_{j})\notin[r_{\beta}]$. Select $h\in H_{r_{\beta}}$ such that $h$ substitutes all variables in $\beta$ to some letter, i.e., $|h(x)|=1$ for all $x\in\operatorname{var}(\beta)$. In particular, let $h(x_{i})=h(x_{k})=\mathtt{a}$, for some $\mathtt{a}\in\Sigma$ and all $(\tilde{r},x_{i},x_{k})\in[r_{\beta}]$. Finally, set $h(x_{j})=h(x_{k^{\prime}})=\mathtt{b}\neq\mathtt{a}=h(x_{i})$, for all other variables $x_{k^{\prime}}\in\operatorname{var}(\beta)$ which are not already set to $h(x_{i})=\mathtt{a}$ and some letter $\mathtt{b}\in\Sigma$ with $\mathtt{a}\neq\mathtt{b}$. Then, we have that $h(\beta)\in L_{NE}(\beta,[r_{\beta}])$ but $h(\beta)\notin L_{NE}(\alpha,[r_{\alpha}])$ as there cannot exist any substitution $h^{\prime}\in H_{r_{\alpha}}$ with $h^{\prime}(x_{i})\neq h^{\prime}(x_{j})$ as $(\tilde{r},x_{i},x_{j})\in[r_{\alpha}]$ and by the assumption that $\tilde{r}$ is an equivalence relation. This is a contradiction to $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$. So, we must have $r_{\alpha}]=[r_{\beta}]$. ∎
Proof of Theorem˜5.1:
Proof
($\Rightarrow$): Assume $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$. As $\tilde{r}$ is an equivalence relation, we get by Lemma 3 that $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$ and $L_{NE}(\beta,r_{\beta})=L_{NE}(\beta,[r_{\beta}])$. Hence, $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$, and we get by Lemma 5 that $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$. Thus, we have $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
($\Leftarrow$): Assume $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$. Then, by Lemma 5 we have $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$ and by Lemma 3 we get that $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ using the same arguments from the other direction. ∎
Proof of Corollary˜3:
Proof
First, we notice that all relations in $\{r_{=},r_{ab},r_{con^{+}}\}$ are equivalence relations that are antisymmetric on elements of length $1$. Now, we normalize $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ such that all variables are introduced with an index in increasing manner, starting with $x_{1}$ in both relational patterns. Then, we check in linear time whether $\alpha=\beta$. If that is the case, we check whether $r_{\alpha}]=[r_{\beta}]$ by constructing the sets representing reflexive, symmetric, and transitive closure of related variables. This can be done in linear time in the lengths of the patterns by scanning them and filling up the resulting sets. The total number of elements in these sets will not exceed the total number of variables in the patterns. If these sets are equal for both relational patterns, we return true, applying Theorem˜5.1. In any other case, we return false. ∎

### v2

@@ 0.B.2 Proof of Lemma˜4
with $(r_{rev},x_{1},x_{1}^{\prime})$, $(r_{rev},x_{2_{1}},x_{2_{2}})$, $(r_{rev},x_{2_{1}},x_{2_{3}})$, $(r_{rev},x_{2_{1}},x_{2_{4}})$, $(r_{rev},x_{2_{1}},x_{2_{5}})$, $(r_{rev},x_{i_{1}},x_{i_{2}})\in r_{\beta_{A}}$ for all $i\in\{3,\ldots,n\}$ and there is a morphism $\sigma:X^{*}\rightarrow\Sigma^{*}$ with $\sigma(\alpha)=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$, then $\sigma(x_{1})=\sigma(x_{1}^{\prime})=a_{1}$ and $\sigma(x_{i_{j}})=a_{i}$ for each $i\in\{1,\ldots,n\}$.
Proof.
@@
must hold. Now, $\sigma(x_{2_{j}})=a_{2}$ must hold for every $j\in\{1,2,3,4,5\}$ since $(r_{rev},x_{2_{1}},x_{2_{2}})$, $(r_{rev},x_{2_{1}},x_{2_{3}})$, $(r_{rev},x_{2_{1}},x_{2_{4}})$, $(r_{rev},x_{2_{1}},x_{2_{5}})$, $(r_{rev},x_{i_{1}},x_{i_{2}})\in r_{\beta_{A}}$ for all $i\in\{3,\ldots,n\}$. Thus, we also get that $\sigma(x_{i_{1}})=\sigma(x_{i_{2}})=a_{i}$ for every $i\in\{3,\ldots,n\}$. ∎
@@
for each $i\in[n]$. Again, all $y_{j,k},\hat{y}_{j,k},\ddot{z}_{j,k}^{\prime}$ are pairwise different variables with $(r_{rev},\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i-1,i}^{\prime}),(r_{rev},\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i,i}^{\prime})\in R$. Now Lemma˜10 applies (mutatis mutandis) as for binary alphabets, and since all substitutions of good form behave for $\Sigma$ as for the binary alphabet, we can use the very same predicates and the same reasoning as before to prove undecidability of the inclusion problem under $r_{rev}$.
This concludes the proof for $r_{rev}$. ∎
0.B.3.3 (The Full Proof Idea for $r_{ab}$)
Proof.
This proof is analogous to the one for $r_{rev}$. Therefore, we only describe how to get to the analogy and refer for a detailed proof to the above part. Let $A$ be a nondeterministic 2-counter automaton. Consider
$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}0$ and
with, for all $i\in\{1,\ldots,\mu\}$, and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{4}},x_{1_{5}},\ldots$ are distinct variables with $(r_{ab},x_{i},x_{i^{\prime}})$, $(r_{ab},x_{i},x_{i_{1}})$, $(r_{ab},x_{i},x_{i_{2}})$, $(x_{i},x_{i_{3}}),(r_{ab},x_{i},x_{i_{4}})$, $(r_{ab},x_{i},x_{i_{5}})\in r_{\beta_{A}}$ and all $\eta_{i}\in X^{*}$ are terminal-free patterns for some $\mu$ and some $\hat{\beta}_{i}$ as for $r_{rev}$. We have $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ with distinct variables $z_{i},z_{i^{\prime}},\hat{z}_{i_{1}},\hat{z}_{i_{2}},\hat{z}_{i_{3}},\hat{z}_{i_{4}},\hat{z}_{i_{5}}$ and $(r_{ab},z_{i},z_{i^{\prime}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{2}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{3}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{4}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in r_{\beta_{A}}$.
The main concept of the proof is to show that iff $L_{E}(\alpha_{A},r_{\alpha_{A}})\subseteq L_{E}(\beta_{a},r_{\beta_{A}})$ then $A$ has no accepting computation. All possible non accepting computations are listed in the $\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}$ part where each $\hat{\beta}_{i},i\in\{1,\ldots,\mu\}$ encodes one non accepting computation together with selecting variables $x_{i_{j}}$ for $j\in{1,2,3}$. The part of $\beta_{A}$ and $\alpha_{A}$ before the first $^{6}$ ensures that exactly one pair $x_{i},x_{i}^{\prime},i\in\{1,\ldots,\mu\}$ is selected, resulting in that exactly one non accepting computation is selected and that exactly one $\ddot{\beta}_{i},i\in\{1,\ldots,\mu\}$ is selected. With this, we get that $\eta_{i},i\in\{1,\ldots,\mu\}$ matches $u$ and thus, that $z_{i}=0$ and $\hat{z}_{i_{1}}=\#$. We use these two variables to encode the 0 and the #\# in the non accepting computation encodings. If we can show that this selecting process also works in the setting $R=r_{ab}$ and that we get two variables with which we can encode the letters 0 and #\#, then this proof also works in this setting. For two variables $x,y$, we have if $x=0$ (or x=#x=\#) and $(x,y)\in r_{ab}$, then $y=0$ (or y=#y=\# respectively). Thus, we can model with the relation $r_{ab}$ identical variables if we know a one letter substitution of one variable.
[…]
$x_{i}$ $x_{i^{\prime}}$ works?
0 0 $00\neq 0\#$
0#0\# 0#0\# or #0\#0 $0#$, $0\#,\ 0\#^{2}0\neq 0\#^{3}$
$0\#^{3}0$ $0\#^{3}0$ $0\#^{3}00\#^{3}0=v\ v$
Table 3: Possibilities for the first match with vvv\ v
Thus, the only possibility is that for exactly one $i\in\{1,\ldots,\mu\}$ the variables $x_{i}$ and $x_{i^{\prime}}$ get substituted with $v$ and all other variables $x_{j},x_{j^{\prime}}$ for $j\neq i,j\in\{1,\ldots,\mu\}$ get substituted with $\epsilon$. This together with the matching from vuvv\ u\ v to $\ddot{\beta}_{1}\ldots\ddot{\beta}_{\mu}$ implies that $\eta_{i}$ needs to match $u$. For this matching between $\eta_{i}=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ and $u=0\#^{5}0$, we get the following possibilities. Remember that $(r_{ab},z_{i},z_{i^{\prime}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{2}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{3}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{4}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in r_{\beta_{A}}$.
$z_{i}$ $\hat{z}_{i_{1}}$ works?
$\epsilon$ 0 or #\# $0^{5},\ \#^{5}\neq 0\#^{5}0$
0 #\# $0\#^{5}0=0\#^{5}0=u$
0#0\# #\# $0\#\#^{5}\#0\neq 0\#^{5}0$
0##0\#\# #\# $0\#^{2}\#^{5}\#^{2}0\neq 0\#^{5}0$
0###0\#\#\# $\epsilon$ $0\#^{6}0\neq 0\#^{5}0$
Table 4: Possibilities for the match with $u$
Thus, the only possibility is that $z_{i}$ and $z_{i^{\prime}}$ get substituted with 0 and $\hat{z}_{i_{k}},k\in\{1,2,3,4,5\}$ get substituted with #\#. With this, we can encode the 0 and the #\# and thus, the proof for $r_{rev}$ works in this setting analogously. ∎
Appendix 0.C Additional Content for the Membership Problem
0.C.1 Proof of Lemma˜5
Proof.
This result follows from known reductions, e.g., [28], for all considered cases in this Lemma. For reference, we recall the construction formally in the relational pattern framework and state why it translates over.
First, we recall positive 1-in-3-SAT. A formula $\varphi$ in positive 3-CNF (conjunctive normal form) is a set of clauses $\varphi=\{C_{1},...,C_{n}\}$, for some $n\in\mathbb{N}$, where each clause $C_{i}$ is a set of 3 non-negated literals $C_{i}=\{X_{i,1},X_{i,2},X_{i,3}\}$, where each $X_{i,j}\in X_{p}$ ($X_{p}$ being the set of all propositional variables). A positive 3-CNF formula $\varphi$ is 1-in-3 satisfiable if there exists an assignment $\beta:X_{p}\rightarrow\{0,1\}$ that sets exactly one literal per clause to $1$.
Assume an arbitrary positive 3-CNF formula $\varphi=\{C_{1},...,C_{n}\}$. Construct the word $w=(1\#)^{n-1}1$ over the binary alphabet $\{1,\#\}$. Next, construct the pattern
$\alpha=x_{1,1}x_{1,2}x_{1,3}\#...\#x_{n,1}x_{n,2}x_{n,3},$
for independent variables $x_{i,j}\in X$ where $i,j\in\mathbb{N}$. Finally, construct the relational constraints $r_{\alpha}$ by adding, for each $X_{i,j}$ and $X_{i^{\prime},j^{\prime}}$ occurring in $\varphi$ where $X_{i,j}=X_{i^{\prime},j^{\prime}}$ are the same variable, the pairs $(x_{i,j},x_{i^{\prime},j^{\prime}}),(x_{i^{\prime},j^{\prime}},x_{i,j})\in r_{\alpha}$. Notice that $|w|_{\#}=|\alpha|_{\#}=n-1$. Hence, each $r_{\alpha}$-valid substitution $h$ must set all variables either to $1$ or to $\varepsilon$ to obtain $h(\alpha)=w$.
Now, we see that if $\varphi$ is 1-in-3-satisfiable, then there exists an assignment of variables $\beta$ that sets exactly one variable per clause to $1$. Based on $\beta$, construct a substitution $h$ where $h(x_{i,j})=1$ iff $\beta{X_{i,j}}=1$. Otherwise set $h(x_{i,j})=\varepsilon$. As each clause only has one variable set to true by $\beta$, we obtain $h(\alpha)=w$. For the other direction, notice that we can analogously construct an assignment $\beta$ out of a $r_{\alpha}$-valid substitution $h$ with $h(\alpha)=w$. This concludes the erasing case. For the non-erasing case, notice that we can construct $(\alpha,r_{\alpha})$ in exactly the same manner and construct a word $w=(1\#)^{4}\#1^{4}$. As the non-erasing case forces each variable to have at least length 1 and all variables must be substituted to multiples of the letter $1$, we can use length $2$ substitutions to represent the positive literals. The argument follows analogously.
  1. 1.
    If $R=\{r_{|w|}\}$, then variables have to be substituted with equal length. As any substitution resulting in $w$ substituted each variable only over unary words over the letter $1$, having equal length also means being the equal word. This concludes this relation for the E and NE cases.
  2. 2.
    If $R=\{r_{ssq}\}$, then relating two variables $x$ and $y$ in both directions, i.e., putting $(r_{ssq},x,y)\in r_{\alpha^{\prime}}$ as well as $(r_{ssq},y,x)\in r_{\alpha^{\prime}}$ results in some substitution $h\in H$ to be valid if and only if $h(x)=h(y)$. This concludes this relation for the E and NE cases.
  3. 3.
    If $R=\{r_{ab}\}$ the same argument as for $r_{|w|}$ holds for the E and NE cases.
  4. 4.
    If $R=\{r_{perm}\}$ the same argument as for $r_{|w|}$ holds for the E and NE cases.
  5. 5.
    If $R=\{r_{rev}\}$ the same argument as for $r_{|w|}$ holds for the E and NE cases.
  6. 6.
    If $R=\{r_{*}\}$ the the same argument as for $r_{ssq}$ holds for the E and NE cases.
  7. 7.
    If $R=\{r_{com^{+}}\}$ then, in the erasing case, substituting one variable to the letter $1$, as $\varepsilon$ is not allowed, forces related variables to also be substituted to the letter $1$ for a $r_{\alpha}$-valid substitution $h$ to result in $h(\alpha)=w$.
  8. 8.
    If $R=\{r_{\varepsilon_{=}}\}$ the sane argument as for $r_{com^{+}}$ holds for the E case.
This concludes this proof. ∎
0.C.2 Proof of Theorem˜5.1
Proof.
Indeed, NP-containment of all considered cases follows from Proposition˜2. For most cases, NP-hardness has been confirmed in 5. What is left to show is that, first, in the case of non-erasing pattern languages, NP-hardness also holds for the relations $r_{com^{+}}$ and $r_{com^{*}}$. Indeed, in the NE case, these two relations result in the same languages, hence they can be considered together. Second, in the erasing case, the relation $r_{com^{*}}$ is still open. The rest of this proof provides two reductions, showing NP-hardness for the remaining two cases.
(Non-Erasing ($r_{com^{+}}$$r_{com^{*}}$): W.l.o.g. we consider $R=\{r_{com^{+}}\}$. We reduce the general 3-SAT problem to the respective membership problem. Let $\varphi=\{C_{1},...,C_{n}\}$ be a 3-CNF formula, for some $n>0$, where $C_{i}=\{X_{i,1},X_{i,2},X_{i,3}\}$ for clauses $C_{i}$ and propositional variables $X_{i,j}\in\{X_{1},...,X_{m},\overline{X_{1}},...,\overline{X_{m}}\}$. Notice that in this variant of the problem, negated variables may occur as literals. First, we construct the word
$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
where $s_{i}=1\#1$ and $t_{j}=1^{10}\#1^{10}\#1^{10}$, for each $1<i\leq n$ and $1\leq j\leq m$. Next, we construct the regular pattern
$\alpha=\#\#\ u_{0}v_{0}\ \#\#\ ...\ \#\#\ u_{m}v_{m}\ \#\#\ \alpha_{1}\ \#\#\ ...\ \#\#\ \alpha_{n}\ \#\#$
where $\alpha_{i}=z_{i,1}\ x_{i,1}\ z_{i,2}\ x_{i,2}\ z_{i,3}\ x_{i,3}\ z_{i,4}$, for $1\leq i\leq n$ and variables $z_{i,1}$ to $z_{i,4}$ as well as $x_{i,1}$ to $x_{i,3}$.
Notice that the pattern $\alpha$ is indeed regular and that each variable occurs only once. Finally, we construct the regular constraints $r_{\alpha}$. For all non-negated variables $X_{k}\in\{X_{1},...,X_{m}\}$ and literals equal to that variable $X_{i,j}=X_{k}$, we add the triple $(r_{com^{*}},u_{k},x_{i,j})\in r_{\alpha}$. Analogously, for all negated variables $\overline{X_{k}}\in\{\overline{X_{1}},...,\overline{X_{m}}\}$, we add the triple $(r_{com^{*}},v_{k},x_{i,j})\in r_{\alpha}$. In conclusion, there exist two parts. First, each factor $s_{k}$, $k\in[m]$, in $w$ has to be matched against $u_{k}v_{k}$ in $\alpha$, and each factor $t_{i}$, $\in[n]$ has to be matched against $\alpha_{i}$. We proceed by proving the arguments.
($\Rightarrow$:) Assume there exists a satisfying assignment of variables $\beta$ for $\varphi$. Let $h$ be a substitution with the following properties: For every variable $X_{k}$, $k\in[m]$, where $\beta(X_{k})=1$ (thus $\beta(\overline{(}X_{k}))=0$), we set $h(u_{k})=1$, $h(v_{k})=\#1$ and, for every literal $X_{i,j}$ with $X_{i,j}=X_{k}$, we set $h(x_{i,j})=1$, and, for every literal with $X_{i,j}=\overline{(}X_{k})$, we set $h(x_{i,j})=\#1$. Analogously, if $\beta(X_{k})=0$ and $\beta(\overline{X_{k}})=1$, we set $h(u_{k})=1\#$ and $h(v_{k})=1$ and the $x_{i,j}$ variables correspondingly. As all variables are assigned some value in $\beta$, all variables of the form $x_{i,j}$ in $\alpha$ are substituted by some value (either $1$ or 1#1\# or #1\#1). For each $\alpha_{i}$, $i\in[n]$, we can now set the surrounding independent variables $z_{i,1}$ to $z_{i,4}$ to non-empty words such that $h(\alpha_{i})=1^{1}0\#1^{1}0\#1^{1}0$. As there is at least one variable $x_{i,j}$ in each $\alpha_{i}$ that is substituted by a word not containing #\# (as $\beta$ is a satisfying assignment), we will not obtain too many #\# by the substitutions for $x_{i,j}$s alone. Additionally, as there are enough $1$s in between each #\# in $w$, we can find a non-empty substitution for each $r_{i,j}$ variable. As all variables related by constraints in $r_{\alpha}$ are substituted by equal words, conforming the relation $r_{com^{+}}$, the chosen substitution is $r_{\alpha}$-valid. We notice that each $h(u_{k}v_{k})=1\#1$, each $h(\alpha_{i})=1^{1}0\#1^{1}0\#1^{1}0$, and therefore $h(\alpha)=w$.
($\Leftarrow$:) Now assume there exists a $r_{\alpha}$-valid substitution $h$ such that $h(\alpha)=w$. We notice that no variable can be substituted to contain the factor ##\#\#, as otherwise $|h(\alpha)|_{\#\#}>|w|_{\#\#}$. Hence, we must have $h(u_{k}vk)=s_{k}=1\#1$ as well as $h(\alpha_{i})=t_{i}=1^{1}0\#1^{1}0\#1^{1}0$. As this is the non-erasing case, for each $k\in[m]$, either $h(u_{k})$ or $h(v_{k})$ contains the symbol #\# while the other is substituted only by $1$. As each $h(\alpha_{i})$ only contains two #\# symbols, there is at least one variable of the form $x_{i,j}$ which is substituted by a unary word in $\{1\}^{+}$. As each such variable is related to some $u_{k}$ or $v_{k}$ respectively, this can only be the case if both are substituted by unary words in $\{1\}^{+}$. Construct an assignment $\beta$ that sets, for each $k\in[m]$, $\beta(X_{k})=1$ and $\beta(\overline{X_{k}})=0$ if $h(u_{k})=1$ (thus $h(v_{k})=\#1$), and, otherwise, set $\beta(X_{k})=0$ and $\beta(\overline{X_{k}})=1$ if $h(v_{k})=1$ (thus $h(u_{k})=1\#$). By the argument from above, we now obtain an assignment that sets at least one variable per clause to $1$, satisfying $\varphi$.
(Erasing ($r_{com^{*}}$): In this case, we can construct a simplified, but analogoue version of the reduction of the NE case. Assume $\varphi$ to be a formula of the form from above. Construct the word
$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
where, this time, $s_{i}=1\#1$ and $t_{j}=1$, for $1\leq i\leq m$ and $1\leq j\leq n$, and let
$\alpha=\#\#\ u_{0}v_{0}\ \#\#\ ...\ \#\#\ u_{m}v_{m}\ \#\#\ \alpha_{1}\ \#\#\ ...\ \#\#\ \alpha_{n}\ \#\#$
where $\alpha_{i}=x_{i,1}x_{i,2}x_{i,3}$ for variables $x_{i,1}$ to $x_{i,3}$. We set $r_{\alpha}$ analogously to the NE case. The first direction of the proof, assuming a satisfying assignment for $\varphi$, we can create a $r_{\alpha}$-substitution similarly to how it is done in the NE case. Set $u_{k}$ and $v_{k}$ analogously to the NE case. For each clause, choose one of the literal variables $x_{i,j}$ that is related to either some $u_{k}$ or $v_{k}$ substituted by $1$ and substitute it by $1$. The other direction of the proof follows by the same reasoning from the NE case that the ##\#\# factors have to align in $w$ and $h(\alpha)$, therefore forcing $h(u_{k}v_{k})=1\#1$ and $h(\alpha_{i})=1$. Hence, for each $i\in[n]$, some variable $x_{i,j}$ is substituted by $1$. This is only possible if either the related $u_{k}$ or $v_{k}$ is substituted by either $\varepsilon$ or $1$ due to the definition of $r_{com^{*}}$. From that, we can construct a satisfying assignment $\beta$ for $\varphi$, analogously to before, concluding this proof idea.
Hence, the membership problem for all cases considered in Theorem˜5.1 is NP-complete, concluding this proof. ∎